A researcher would like to determine whether an over-the-counter cold medication has an effect on mental alertness. A sample of n= 18 participants is obtained, and each person is given a standard dose of the medication one hour before being tested in a driving simulation task. For the general population, scores on the simulation task are normally distributed with μ = 60. The individuals in the sample had an average score of M = 54 and an s = 8.

Can the researcher conclude that scores on the driving simulation task are significantly different after taking the medication? Use a two-tailed test with α = .05. Be sure to state your null hypothesis, a decision about that null hypothesis, and a conclusion explaining your results with an APA format statement. Calculate Cohen’s d if necessary.
Answer #1

Given that, sample size ( n ) = 18

sample mean ( M ) = 54 and

sample standard deviation ( s ) = 8

significance level (alpha) = 0.05

The null and the alternative hypotheses are,

Ho : μ = 60

Ha : μ 60

Test statistic is,

M_μ 54-60一一3.182 7t v/ 18

Test statistic is, t = -3.182

t-critical values at a-0.05 with degrees of freedom = n - 1= 18 -1 = 17 are, t^*= pm 2.110

Here, test statiatic = -3.182 < -2.110

so, we reject the null hypothesis ( H0).

p-value = 0.0055

Conclusion: There is sufficient evidence to conclude that scores on the driving simulation task are significantly different after taking the medication.